OpenStudy (anonymous):

fWhat is the area of the trapezoid if DC = 4x and AB = 2x and the height is (3x + 1)?

jimthompson5910 (jim_thompson5910):

I'm assuming that AB and CD are the bases. If so, then use the formula A = h*(b1+b2)/2 and plug in the two bases to get A = h*(4x+2x)/2 A = h*(6x)/2 A = h*3x then plug in the height 3x+1 to get A = (3x+1)*3x A = 3x(3x+1) A = 3x*3x+3x*1 A = 9x^2 + 3x